# MOTGNN: Interpretable Graph Neural Networks for Multi-Omics Disease Classification

### v1

@@ 3.3. Graph Neural Network and Feature Fusion
We utilize the graph-embedded deep feedforward network (GEDFN) model (Kong and Yu, 2018) as the core GNN component used in MOTGNN. Unlike standard fully connected architectures, GEDFN directly incorporates the adjacency matrix into the weight connections between the input and the first hidden layers. This design enables the model to respect the feature-level relationships encoded in the graph while enforcing the sparsity, a common assumption for omics-based feature graphs. The number of neurons in the first hidden layer must be the same as the input dimension. Formally, given the reduced input data $X^{*}\in\mathbb{R}^{n\times p^{*}}$ and its graph adjacency matrix $\tilde{A}\in\mathbb{R}^{p^{*}\times p^{*}}$ (augmented with self-loops), the first hidden layer $Z_{1}\in\mathbb{R}^{n\times p^{*}}$ is computed as
$Z_{1}=\sigma(X^{*}(W_{in}\odot\tilde{A})+b_{in}),$
where $W_{in}\in\mathbb{R}^{p^{*}\times p^{*}}$ and $b_{in}\in\mathbb{R}^{n\times p^{*}}$ are the weight and bias matrices, $\odot$ denotes element-wise (Hadamard) multiplication, and $\sigma(\cdot)$ is the activation function. The design of GEDFN shares conceptual similarities with the graph attention network (GAT) formulation discussed in (Chen et al., 2023), where the GAT layer is represented as
$Z_{1}=\sigma((W_{attention}\odot A)X+b_{in}),$
with $W_{attention}\in\mathbb{R}^{n\times n}$ and the nodes representing samples rather than features. In this formulation, GAT dynamically learns attention weights over neighbors. Similarly, GEDFN enables dynamic weight learning while ensuring that connections between features and the first hidden layer are strictly supported by the feature graph. This yields a sparse and interpretable architecture tailored to omics data.
Following the training of three omics-specific GNNs, each on its respective dimension-reduced input and constructed graph, we obtain three last hidden representations: $Z_{1}$, $Z_{2}$, and $Z_{3}$. These embeddings are concatenated to form a unified representation:
@@
where $Z$ denotes the concatenated neurons, and is then passed to a DFN, which consists of fully connected layers and a final softmax output to perform binary disease classification.
Table 2. Data description of three TCGA cancer datasets (COADREAD, LGG, and STAD) regarding sample sizes, class distributions, the number of original features, and the number of preprocessed features for each omics type (methylation, mRNA, and miRNA).
@@
Note: meth = DNA methylation, mRNA = gene expression, miRNA = microRNA expression. The processed data were obtained from the Broad GDAC Firehose (Lu et al., 2023).
Figure 2. The histogram of methylation, mRNA, and miRNA on preprocessed COADREAD, LGG, and STAD datasets. Each omics dataset was individually scaled to [0, 1]. Different omics types demonstrate distinct distribution patterns.
[…]
[Omics distribution]The histograms of methylation, mRNA, and miRNA on COADREAD, LGG, and STAD datasets. Each omics dataset was individually scaled to the [0, 1] range using min-max normalization. Different omics data demonstrate distinct distribution patterns.
@@ 3.4. Interpretability on Features and Graphs
To interpret model predictions and evaluate the contribution of features and graphs, we adopt and extend the connection weights method (Olden and Jackson, 2002). Feature importance is assessed by summing the absolute weights connecting each input neuron to the first hidden layer, while accounting for adjacency-based sparsity. For feature $j$ in $X_{i}$, the importance score $IF_{j}^{i}$ is defined as:
$IF_{j}^{i}=\sum_{u=1}^{p_{i}}\left|W_{ju}^{(in)}I(\tilde{A}_{ju}^{i}=1)\right|,$
where $W^{(\text{in})}$ is the input weight matrix and $I(\cdot)$ is the indicator function. After training the model, each input feature is assigned an importance score. These scores can be used to rank the features, allowing the selection of top biomarkers for further biological investigation. To evaluate the relative importance of the graph for each omics modality, we compute the $L_{1}$ norm of the weight matrix connecting the GNN’s last hidden representations ($Z_{i}$) to the DFN layer. The relative graph importance (RIG) is then:
@@
where $i=1,2,3$, $W_{Z_{i}\leftrightarrow f}$ is the weight matrix connecting $Z_{i}$ and DFN, and $\|\cdot\|_{1}$ is the $L_{1}$ norm that sums the absolute values of its inside matrix. The total contribution sums to one across all three graphs.
@@ 3.5. Technical Details
We implement the xgboost Python library (Chen and Guestrin, 2016) for the XGBoost module in MOTGNN. We tuned the number of estimators (determine trees). We employ the Tensorflow library (Abadi et al., 2016) for training GNN and DFN modules. Model optimization uses the Adam optimizer (Kingma and Ba, 2014) with mini-batch training. The nonlinear activation function uses the Rectified Linear Unit (ReLU) function: $\sigma(x)=max(0,x)$ (Nair and Hinton, 2010). For classification, the final output layer uses a softmax function, and the loss is computed using binary cross-entropy:
$L(Y,\hat{Y})=-\frac{1}{n}\sum_{i=1}^{n}\left\{y_{i}\ln(\hat{y}_{i})+(1-y_{i})\ln(1-\hat{y}_{i})\right\}.$
We tune hyperparameters across three main categories: (1) architectural (number of layers and hidden neurons), (2) training-related (learning rate, batch size, and epochs), and (3) regularization-related (dropout and L2 penalty). To avoid overfitting, we apply strategies including dropout, batch normalization, L2 regularization, and early stopping. Data are split into training, validation, and test sets in a 60%:20%:20% ratio, with stratified sampling to address class imbalance. We have conducted 20 independent and reproducible train-validation-test splits. Hyperparameters are tuned via grid search in a reasonable space and based on experience.
@@ 4.1. Data Description
We evaluate our method on three real-world cancer datasets from The Cancer Genome Atlas (TCGA) (Weinstein et al., 2013): COADREAD, STAD, and LGG. The processed multi-omics data and label annotations were obtained from the Broad GDAC Firehose (https://gdac.broadinstitute.org/). Each dataset includes matched DNA methylation ($X_{1}$), mRNA expression ($X_{2}$), and miRNA expression ($X_{3}$) data. Only samples with complete data across all three omics types were included in our analysis. The COADREAD dataset combines COlorectal ADenocarcinoma (COAD) and REctal ADenocarcinoma (READ) into a binary classification task, with 254 COAD and 78 READ samples. The LGG dataset involves grade classification for low-grade gliomas, with 255 Grade 2 and 269 Grade 3 samples. The STAD dataset targets the classification of STomach ADenocarcinoma subtypes, with 205 samples from adenocarcinoma (ADC) and 166 from intestinal adenocarcinoma (IAC). Among the three datasets, COADREAD has the most imbalanced class distribution.
Table 2 summarizes the datasets regarding sample sizes, class distributions, the number of original features, and the number of preprocessed features for each omics type. Data preprocessing includes quality control steps to remove noisy and redundant features, such as filtering out zero-expression signals, low-variance features, and handling multiple comparisons (Lu et al., 2023). Each omics dataset was independently normalized to the range [0, 1] using min-max scaling. The preprocessed multi-omics data shared in (Lu et al., 2023) contains minor inconsistencies in sample sizes across omics modalities. For the COADREAD dataset, meth and mRNA data include 332 samples, while miRNA contains 337 samples. We addressed this discrepancy by removing the five extra miRNA samples to ensure alignment. Similarly, in the STAD dataset, meth and mRNA data comprise 371 samples, whereas miRNA includes 372 samples. Here, we excluded one redundant miRNA sample to maintain consistency.
The histograms in Figure 2 show the distribution of feature values across three omics types (methylation, mRNA, and miRNA) for the COADREAD, LGG, and STAD datasets. The distinct distributional patterns observed across omics types within each dataset highlight each modality’s inherent heterogeneity and diverse data characteristics. These differences across the three datasets help evaluate the potential of MOTGNN’s generalizability.
@@ 4.2. Comparison and Evaluation Metrics
We evaluate the performance of our proposed model, MOTGNN, in comparison to several baseline machine learning and deep learning models: XGBoost (XGB) (Chen and Guestrin, 2016), random forest (RF) (Breiman, 2001), deep feedforward network (DFN) (Goodfellow et al., 2016), and graph convolutional network (GCN) (Kipf and Welling, 2017). The number of trees for XGBoost and RF is both at the default, which is 100. The number of hidden layers for the DFN and GCN models is two, with several hidden neurons of 64 per layer. All baseline models input the column-wise concatenation of the three preprocessed omics data types (methylation, mRNA, and miRNA). Table 3 summarizes the tuning ranges and selected values (optimal choices in bold). We report three standard evaluation metrics: accuracy, ROC-AUC, and F1-score. Accuracy measures the proportion of correctly classified samples. ROC-AUC (area under the receiver operating characteristic curve) (Green et al., 1966) quantifies the model’s ability to distinguish between classes. The F1-score is defined as $F1=\frac{2\times\text{precision}\times\text{recall}}{\text{precision}+\text{recall}},$ where precision refers to the positive predictive value, and recall corresponds to sensitivity or true positive rate. As the harmonic mean of precision and recall, the F1-score provides a balanced evaluation, especially in scenarios with imbalanced class distributions, and mitigates bias from simply predicting the majority class. These metrics comprehensively assess classification performance across different datasets and modeling approaches. We conducted all computations on the Falcon supercomputer (Idaho C3+3 Collaboration, 2022), requesting four CPUs and 20 GB of memory per job.
Table 3. Hyperparameter tuning for XGBoost and neural networks components in MOTGNN on COADREAD, LGG, and STAD datasets. The optimal choices are in bold.
@@
Figure 3. Feature dimensions before and after XGBoost on COADREAD, LGG, and STAD datasets. The bar plots compare the preprocessed feature dimensions (pre-selection) against the reduced dimensions (post-selection).
[…]
[Dimension reduction]Feature Dimension before and after XGBoost selection
5. Results and Discussion
Table 4. Structural properties of sparse graphs constructed by MOTGNN for COADREAD, LGG, and STAD datasets.
@@ 5. Results and Discussion
Note: The number of edges includes the self-loops, $E_{i}$ is the set of edges for Graph $G_{i}$, $m_{i}$ is the edge to node ratio with $m_{i}=\frac{E_{i}}{p_{i}^{*}}$.
Figure 4. Relative graph importance of different omics types (methylation, mRNA, and miRNA) in MOTGNN’s classification performance across COADREAD, LGG, and STAD datasets. Bar heights represent each omics layer’s contribution, revealing dataset-specific biological dominance.
[…]
[Relative graph importance]Relative Graph Importance of different omics type (meth, mRNA, and miRNA) for Classification Task on COADREAD, LGG, and STAD datasets
Note: $G_{i}$ = Graphs from reduced features (methylation:mRNA:miRNA), $RIG_{i}$ = Relative importance of graph $G_{i}$ in MOTGNN (normalized to sum to 1.0 per dataset).
Table 5. Classification performance comparison on COADREAD, LGG, and STAD datasets across different evaluation metrics (Mean ± SD [95% CI]). MOTGNN consistently outperforms baseline models across all datasets, highlighting its robustness in integrating multi-omics data (Best results are bolded)
@@
GCN 0.871 ± 0.051 [0.847, 0.895] 0.92 ± 0.049 [0.897, 0.942] 0.735 ± 0.107 [0.685, 0.785]
@@
GCN 0.692 ± 0.051 [0.668, 0.716] 0.763 ± 0.049 [0.74, 0.785] 0.693 ± 0.057 [0.666, 0.719]
@@
GCN 0.64 ± 0.047 [0.618, 0.662] 0.694 ± 0.045 [0.673, 0.715] 0.603 ± 0.053 [0.578, 0.628]
@@
Note: Results are reported as mean ± standard deviation [95% confidence interval] over 20 runs (n=20). Confidence intervals are calculated using the t-distribution: $\text{CI}=\bar{x}\pm t_{0.975,19}\times\frac{SD}{\sqrt{n}}$ where $n=20$, $t_{0.975,19}=2.093$. All intervals are symmetric about the mean.
@@
Figure 5. Classification performance comparison for the COADREAD, LGG, and STAD datasets. Bars show mean scores; error bars represent 95% confidence intervals over 20 independent runs.
[…]
[Classification Plot]Classification Bar Plots with Error Bars
Figure 3 presents the feature dimensions before and after applying XGBoost within the MOTGNN framework. The reduced omics matrices $X_{1}^{*}$, $X_{2}^{*}$, and $X_{3}^{*}$ represent the selected feature subsets for DNA methylation, mRNA, and miRNA, respectively. Note that the reduced feature sizes ($p_{i}^{*}$) vary across omics types due to the nature of XGBoost’s feature selection: the final selected set is the union of features used across all $M$ decision trees (see details in Section 3.2). Table 4 highlights the structural properties across MOTGNN’s constructed feature graphs ($G_{1}$-$G_{3}$). Figure 4 presents the relative graph importance for the three constructed graphs ($G_{1}$-$G_{3}$). The edge-to-node ratios (2.1-2.8) suggest sparse graph structures, with miRNA ($G_{3}$) exhibiting the highest edge density (2.35-2.79) but the lowest relative importance (24-29%), indicating limited discriminative value. In COADREAD, mRNA ($G_{2}$) dominates (42% importance), while methylation ($G_{1}$) and mRNA ($G_{2}$) contribute equally in LGG and STAD (35-37%). The consistently high methylation importance (34-37%) across datasets underscores its robustness as a biomarker, whereas mRNA’s variability reflects tissue-specific expression dynamics. These results demonstrate MOTGNN’s ability to prioritize informative omics modalities while maintaining computational efficiency through sparse graph structures, with edge counts remaining tractable for clinical-scale applications.
Table 5 presents classification results averaged over 20 independent train-validation-test splits. Our proposed MOTGNN model consistently outperforms all baselines (XGBoost, RF, DFN, and GCN) across all metrics (Accuracy, ROC-AUC, and F1-score) on all three datasets (COADREAD, LGG, and STAD). In COADREAD, MOTGNN achieves an accuracy of 93.9%, outperforming XGBoost (91.0%), RF (80.1%), DFN (87.3%), and GCN (87.1%). The performance gap is even more evident in F1-score: MOTGNN attains 87.2%, marking a 9.3% absolute improvement over XGBoost (77.9%) and a substantial 53.8% advantage over RF (33.4%). Given COADREAD’s imbalanced class distribution (254:78), this improvement highlights MOTGNN’s robustness in handling skewed datasets, where traditional models like RF show substantially reduced performance. The ROC-AUC results further validate MOTGNN’s stability, with scores exceeding 96.9% for COADREAD and maintaining at least 3% improvements over the best baseline across all datasets. While XGBoost performs competitively in COADREAD (96.4%), its inconsistency becomes apparent when examining results across datasets, dropping to 71.8% for LGG and 63.5% for STAD. DFN shows stronger baseline performance, particularly in STAD (69.1% ROC-AUC), but still falls short of MOTGNN’s 71.2%. GCN exhibits moderate performance across datasets and outperforms DFN slightly in LGG and STAD, which suggests some benefit from modeling feature relationships through a graph structure. However, GCN constructs only one graph over concatenated multi-omics features, which may weaken omics-specific signals and introduce noise from irrelevant cross-omics interactions. In contrast, MOTGNN’s modality-specific graph-based architecture provides additional performance gains, particularly under class imbalance, highlighting its suitability for complex biomedical classification tasks.
@@
Figure 6. F1-score comparison across 20 runs on the imbalanced COADREAD dataset. (A) Box plots show median (labeled), interquartile range, and outliers. (B) Violin plots show median (labeled) and illustrate score distributions. MOTGNN demonstrates both high performance and low variability.
[…]
[BoxandViolinPlot]Box and violin plots showing F1-score distributions for different models on COADREAD.

### v2

@@ 2.3. Graph Neural Network and Feature Fusion
We use the graph-embedded deep feedforward network (GEDFN) model (Kong and Yu, 2018) as the core GNN component used in MOTGNN. Unlike standard fully connected architectures, GEDFN directly incorporates the adjacency matrix into the weight connections between the input and the first hidden layer. This design allows the model to respect the feature-level relationships encoded in the graph while enforcing sparsity, which is a common assumption for omics-based feature graphs. The number of neurons in the first hidden layer is set equal to the input dimension. Given a reduced input data $X^{*}\in\mathbb{R}^{n\times p^{*}}$ and its augmented graph adjacency matrix $\tilde{A}\in\mathbb{R}^{p^{*}\times p^{*}}$, the first hidden layer $Z_{1}\in\mathbb{R}^{n\times p^{*}}$ is computed as
$Z_{1}=\sigma(X^{*}(W_{\text{in}}\odot\tilde{A})+b_{\text{in}}),$
where $W_{\text{in}}\in\mathbb{R}^{p^{*}\times p^{*}}$ and $b_{\text{in}}\in\mathbb{R}^{n\times p^{*}}$ are the weight and bias matrices, $\odot$ denotes element-wise (Hadamard) multiplication, and $\sigma(\cdot)$ is the activation function. This ensures that only graph-supported connections are active between the input features and the first hidden layer. The design of GEDFN shares conceptual similarities with the graph attention network (GAT) formulation discussed in (Chen et al., 2023), where the GAT layer is represented as
$Z_{1}=\sigma((W_{\text{attention}}\odot A)X+b_{\text{in}}),$
with $W_{\text{attention}}\in\mathbb{R}^{n\times n}$ and the nodes representing samples rather than features. In this formulation, GAT dynamically learns attention weights over neighbors. In GEDFN, the graph structure constrains which connections are allowed, while the model still learns connection strengths. This yields a sparse and interpretable architecture tailored to omics data.
After training three omics-specific GNNs, each on its own dimension-reduced input and constructed graph, we obtain three final hidden representations: $Z_{1}$, $Z_{2}$, and $Z_{3}$. These embeddings are concatenated to form a unified representation:
@@
which is then passed to a DFN consisting of fully connected layers and a final softmax output layer to perform binary disease classification.
Table 2. Summary of three TCGA cancer datasets (COADREAD, LGG, and STAD), including sample sizes, class distributions, the number of original features, and the number of preprocessed features for each omics type: DNA methylation (meth), gene expression (mRNA), and microRNA expression (miRNA). The preprocessed data were obtained from the Broad GDAC Firehose.
@@
Figure 2. Distribution of DNA methylation, mRNA, and miRNA features across the preprocessed COADREAD, LGG, and STAD datasets. Each omics dataset was independently scaled to the range of [0, 1] using min-max normalization. The distinct distributional patterns highlight the heterogeneous characteristics of different omics types.
@@ 2.4. Interpretability on Features and Graphs
To interpret model predictions and evaluate the contribution of both features and graphs, we adopt and extend the connection weights method (Olden and Jackson, 2002). Feature importance is quantified by summing the absolute values of weights connecting each input neuron to the first hidden layer, while accounting for adjacency-based sparsity in the feature graph. For feature $j$ in omics modality $X_{i}$, the importance score $IF_{j}^{i}$ is defined as:
$IF_{j}^{i}=\sum_{u=1}^{p_{i}}\left|W_{ju}^{(\text{in})}I(\tilde{A}_{ju}^{i}=1)\right|,$
where $i=1,2,3$, $W^{(\text{in})}$ is the input weight matrix and $I(\cdot)$ is the indicator function. The final importance score for feature $j$ across all omics modalities is obtained by summing its importance values:
$IF_{j}=IF_{j}^{1}+IF_{j}^{2}+IF_{j}^{3}.$
After training the model, each input feature is assigned an importance score that can be used to rank and identify top biomarkers for downstream biological investigation. To assess the relative importance of each omics-specific graph, we compute the $L_{1}$ norm of the weight matrix connecting the GNN’s last hidden representations ($Z_{i}$) to the DFN layer. The relative graph importance ($RIG_{i}$) is defined as:
@@
where $W_{Z_{i}\leftrightarrow f}$ denotes the weight matrix connecting $Z_{i}$ to the DFN, and $\|\cdot\|_{1}$ is the $L_{1}$ norm (the sum of absolute values). The total relative importance across all three graphs sums to one.
@@ 2.5. Technical Details
We implement the XGBoost module in MOTGNN using the xgboost Python library and tune the number of estimators (i.e., the number of trees). The GNN and DFN modules are implemented in the Tensorflow library (Abadi et al., 2016). Model optimization uses the Adam optimizer (Kingma and Ba, 2014) with mini-batch training. The nonlinear activation function uses the Rectified Linear Unit (ReLU) function: $\sigma(x)=max(0,x)$ (Nair and Hinton, 2010). For classification, the final output layer uses a softmax function, and the loss is computed using binary cross-entropy:
$L(Y,\hat{Y})=-\frac{1}{n}\sum_{i=1}^{n}\left[y_{i}\ln(\hat{y}_{i})+(1-y_{i})\ln(1-\hat{y}_{i})\right],$
where $y_{i}\in\{0,1\}$ is the ground truth label, and $\hat{y}_{i}$ is the predicted probability of the positive class.
We tune hyperparameters across three main categories: (1) architectural (number of layers and hidden neurons), (2) training-related (learning rate, batch size, and epochs), and (3) regularization-related (dropout and $L_{2}$ penalty). The $L_{2}$ regularization term is expressed as $L_{\text{reg}}=\lambda\sum_{j}w_{j}^{2}$, where $\lambda$ is the regularization coefficient and $w_{j}$ denotes model weights. To avoid overfitting, we apply various strategies including dropout, batch normalization, $L_{2}$ regularization, and early stopping. Data are split into training, validation, and test sets in a 60%:20%:20% ratio, with stratified sampling used to address class imbalance. We perform 20 independent and reproducible train-validation-test splits. Hyperparameters are tuned via a grid search within a reasonable space and informed by prior experience.
@@ 3.1. Data Description
We evaluate our method on three real-world cancer datasets from The Cancer Genome Atlas (TCGA) (Weinstein et al., 2013): COADREAD, STAD, and LGG. The processed multi-omics data and label annotations were obtained from the Broad GDAC Firehose (https://gdac.broadinstitute.org/). Each dataset includes matched DNA methylation ($X_{1}$), mRNA expression ($X_{2}$), and miRNA expression ($X_{3}$), along with class labels ($Y$). Only samples with complete measurements across all three omics types were included in our analysis. The COADREAD dataset combines COlorectal ADenocarcinoma (COAD) and REctal ADenocarcinoma (READ) into a binary classification task, with 254 COAD and 78 READ samples. The LGG dataset involves grade classification for low-grade gliomas, with 255 Grade 2 and 269 Grade 3 samples. The STAD dataset targets the classification of STomach ADenocarcinoma subtypes, with 205 adenocarcinoma (ADC) and 166 intestinal adenocarcinoma (IAC) samples. Among the three datasets, COADREAD exhibits the most imbalanced class distribution.
Table 2 summarizes the datasets in terms of sample sizes, class distributions, and the number of original and preprocessed features for each omics type. Data preprocessing included standard quality control steps to remove noisy and redundant features, such as filtering out zero-expression signals and low-variance features, as well as adjusting for multiple comparisons (Lu et al., 2023). Each omics dataset was independently normalized to the range [0, 1] using min-max scaling. The preprocessed data shared in (Lu et al., 2023) contained minor inconsistencies in sample sizes across omics types. For COADREAD, methylation and mRNA data each included 332 samples, while miRNA had 337. We removed the five extra miRNA samples to ensure alignment. Similarly, for STAD, methylation and mRNA data each had 371 samples, whereas miRNA had 372; we excluded one redundant miRNA sample for consistency. The histograms in Figure 2 show the feature distributions across three omics types (methylation, mRNA, and miRNA) for COADREAD, LGG, and STAD. The distinct patterns among omics types highlight the inherent heterogeneity and complementary nature of each modality. These differences across datasets help evaluate the generalizability of MOTGNN.
@@ 3.2. Comparison and Evaluation Metrics
We evaluate the performance of our proposed model, MOTGNN, which by default employs XGBoost to generate three graphs (one per omics type) and uses GEDFN as the main graph neural network (GNN) component. The model is compared with four baseline machine learning and deep learning models: XGBoost, random forest (RF), deep feedforward network (DFN), and graph convolutional network (GCN). To assess the contribution of individual architectural components, we also design two ablation variants of MOTGNN: (1) $\mathrm{MOTGNN}_{\text{gcn}}$ replaces the GEDFN module with a standard GCN, and (2) $\mathrm{MOTGNN}_{\text{rf}}$ uses RF instead of XGBoost for graph generation. These ablation models share the same overall architecture as MOTGNN but differ only in the modified component.
The number of trees for both XGBoost and RF is set to 100. The DFN and GCN models each have two hidden layers with 64 neurons per layer. For XGBoost, RF, DFN, and GCN, the input consists of the column-wise concatenation of the three preprocessed omics data types (methylation, mRNA, and miRNA). In contrast, $\mathrm{MOTGNN}_{\text{gcn}}$ and $\mathrm{MOTGNN}_{\text{rf}}$ process each omics modality separately, using three parallel GNNs, one for each omics type. The graphs used in GCN and $\mathrm{MOTGNN}_{\text{gcn}}$ are generated by XGBoost, while $\mathrm{MOTGNN}_{\text{rf}}$ are generated by RF. Table 3 summarizes the hyperparameter tuning ranges and selected values (optimal choices in bold) of MOTGNN.
We report three standard evaluation metrics: accuracy, ROC-AUC, and F1-score. Accuracy measures the proportion of correctly classified samples. ROC-AUC (area under the receiver operating characteristic curve) quantifies the model’s ability to distinguish between classes. F1-score is defined as
$F1=\frac{2\times\text{precision}\times\text{recall}}{\text{precision}+\text{recall}},$
where precision refers to the positive predictive value and recall corresponds to sensitivity (true positive rate). As the harmonic mean of precision and recall, the F1-score provides a balanced assessment under imbalanced class distributions, mitigating bias toward the majority class. All computations were conducted on the Falcon supercomputer (Idaho C3+3 Collaboration, 2022), using four CPUs and 20 GB of memory per job.
Table 3. Hyperparameter tuning ranges for the XGBoost and neural network components in MOTGNN across the COADREAD, LGG, and STAD datasets. Optimal values are highlighted in bold.
@@
Figure 3. Feature dimensions before and after XGBoost-based selection on COADREAD, LGG, and STAD datasets. The bar plots compare preprocessed feature dimensions (pre-selection) with reduced dimensions (post-selection), showing substantial dimensionality reduction across all omics types.
Table 4. Structural properties of sparse graphs constructed by MOTGNN for COADREAD, LGG, and STAD datasets. $E_{i}$ denotes the number of edges (including self-loops) in graph $G_{i}$, and $m_{i}$ represents the edge-to-node ratio, $m_{i}=E_{i}/p_{i}^{*}$.
@@
Table 5. Classification performance comparison of MOTGNN and other models on COADREAD, LGG, and STAD datasets. Results are reported as mean ± standard deviation [95% confidence interval] over 20 independent runs. Baseline models include XGBoost, RF (random forest), DFN (deep feedforward network), and GCN (graph convolutional network). $\mathrm{MOTGNN}_{\text{gcn}}$ and $\mathrm{MOTGNN}_{\text{rf}}$ represent MOTGNN variants that respectively replace GEDFN with GCN or use RF instead of XGBoost for graph generation. Bold values indicate the best performance.
@@
GCN 0.871 ± 0.051 [0.847, 0.895] 0.920 ± 0.049 [0.897, 0.942] 0.735 ± 0.107 [0.685, 0.785]
$\mathrm{MOTGNN}_{\text{gcn}}$ 0.895 ± 0.039 [0.877, 0.913] 0.928 ± 0.034 [0.912, 0.944] 0.783 ± 0.078 [0.746, 0.820]
$\mathrm{MOTGNN}_{\text{rf}}$ 0.933 ± 0.028 [0.920, 0.946] 0.967 ± 0.026 [0.955, 0.979] 0.859 ± 0.061 [0.830, 0.888]
@@
GCN 0.692 ± 0.051 [0.668, 0.716] 0.763 ± 0.049 [0.740, 0.785] 0.693 ± 0.057 [0.666, 0.719]
$\mathrm{MOTGNN}_{\text{gcn}}$ 0.696 ± 0.050 [0.673, 0.719] 0.764 ± 0.047 [0.742, 0.786] 0.692 ± 0.052 [0.668, 0.716]
$\mathrm{MOTGNN}_{\text{rf}}$ 0.706 ± 0.030 [0.692, 0.720] 0.777 ± 0.038 [0.759, 0.795] 0.703 ± 0.036 [0.686, 0.720]
@@
GCN 0.640 ± 0.047 [0.618, 0.662] 0.694 ± 0.045 [0.673, 0.715] 0.603 ± 0.053 [0.578, 0.628]
$\mathrm{MOTGNN}_{\text{gcn}}$ 0.653 ± 0.062 [0.624, 0.682] 0.686 ± 0.057 [0.659, 0.713] 0.613 ± 0.076 [0.577, 0.649]
$\mathrm{MOTGNN}_{\text{rf}}$ 0.657 ± 0.048 [0.635, 0.679] 0.710 ± 0.049 [0.687, 0.733] 0.616 ± 0.058 [0.589, 0.643]
@@
Figure 4. Comparison of classification performance across COADREAD, LGG, and STAD datasets. Bars show mean scores, and error bars represent 95% confidence intervals over 20 independent runs. MOTGNN consistently achieves the highest scores across accuracy, ROC-AUC, and F1 metrics.
@@
Figure 5. F1-score comparison across 20 independent runs on the imbalanced COADREAD dataset (class ratio 254:78). (A) Box plots display the median (labeled), interquartile range, and outliers. (B) Violin plots illustrate the distribution of F1-scores with labeled medians. MOTGNN achieves the highest median score with the lowest variability, demonstrating stable and robust performance under class imbalance.
Figure 6. Relative importance of feature graphs from different omics types (methylation, mRNA, and miRNA) in MOTGNN across COADREAD, LGG, and STAD datasets. Bar heights indicate the contribution of each omics layer to the overall classification performance, revealing dataset-specific biological dominance patterns.
4. Results and Discussion
Feature Selection and Graph Properties. Figure 3 presents the feature dimensions before and after applying XGBoost within the MOTGNN framework. The reduced omics matrices $X_{1}^{*}$, $X_{2}^{*}$, and $X_{3}^{*}$ represent the selected feature subsets for DNA methylation, mRNA, and miRNA, respectively. The reduced feature sizes ($p_{i}^{*}$) vary across omics types due to the nature of XGBoost’s supervised feature selection: the final selected set is the union of features used across all $M$ decision trees, as described in Section 2.2. Table 4 summarizes the structural properties of the three constructed feature graphs ($G_{1}$, $G_{2}$, $G_{3}$), where $E_{i}$ denotes the set of edges for Graph $G_{i}$ and the edge-to-node ratio is defined as $m_{i}=\frac{E_{i}}{p_{i}^{*}}$ for $i=1,2,3$.
Classification Performance. Table 5 presents the classification results averaged over 20 independent train-validation-test splits. Overall, MOTGNN consistently outperforms all baseline models (XGBoost, RF, DFN, and GCN) across all evaluation metrics (accuracy, ROC-AUC, and F1-score) on all datasets (COADREAD, LGG, and STAD). In COADREAD, MOTGNN achieves an accuracy of 93.9%, surpassing XGBoost (91.0%), RF (80.1%), DFN (87.3%), and GCN (87.1%). The improvement is even more pronounced in F1-score, where MOTGNN attains 87.2%, marking a 9.3% absolute improvement over XGBoost (77.9%) and a substantial 53.8% advantage over RF (33.4%). Given COADREAD’s imbalanced class distribution (254:78), these gains highlight MOTGNN’s robustness under skewed data conditions, where traditional ensemble models degrade substantially. The two ablation variants, $\mathrm{MOTGNN}_{\text{gcn}}$ and $\mathrm{MOTGNN}_{\text{rf}}$, both outperform their respective base counterparts (GCN and RF), confirming the benefits of the modality-specific multi-graph design. However, the full MOTGNN remains superior, demonstrating that combining XGBoost-based graph construction with the GEDFN module provides complementary advantages for capturing nonlinear and omics-specific dependencies.
Across all datasets, MOTGNN maintains the highest ROC-AUC values ($\geq$ 96.9% for COADREAD and $\geq$ 3% higher than the best baseline in LGG and STAD). While XGBoost performs well in COADREAD (96.4%), its performance declines on LGG (71.8%) and STAD (63.5%), indicating limited generalizability. DFN exhibits competitive baseline performance, particularly in STAD (69.1% ROC-AUC), but lacks the ability to capture inter-feature relationships effectively. GCN achieves moderate gains by leveraging graph structure, but it constructs a single concatenated graph across omics, which may weaken omics-specific signals and introduce noise from irrelevant cross-omics interactions. By contrast, MOTGNN learns individual omics-specific graphs and integrates them through interpretable GNN modules, achieving consistent and robust performance across all datasets. These results confirm its effectiveness for complex multi-omics classification tasks.